What is 25% off of $527
Answer by VirtualMathTutor(26)

Find 25% of $527, then subtract the answer from $527
25% of $527 is 0.25 * 527 = $131.75

So 25% off of $527 will be $527 - $131.75 = $395.25


Hope this helps.


Another approach is: 100% - 25% = 75%


So we will find 75% of 527 = 0.75 * 527 = $395.25


You will get the same answer as the first approach.